OpenStudy (anonymous):

The volume of a solid sphere of radius r is given by the equation V=4/3TTr^3. Derive this equation by using either the disk or shell method for finding the volume of a solid of revolution.

OpenStudy (anonymous):

a circle is x^2 + y^2 = r^2 , we can revolve the region bounded by y=0 and y = sqrt (r^2 - x^2) about the x axis using disc method

OpenStudy (anonymous):

with limits of integration from -r to r

OpenStudy (anonymous):

so integral pi * [ sqrt (r^2 - x^2) ] ^2 from -r to r

OpenStudy (anonymous):

quite easy..i am in sixth grade and know that ...how old are u The volume of a sphere is (4/3)*pi*r3 So if you are given a volume the radius is r = {Vol / ((4/3) *(pi))}(1/3) The 1/3 means the cubic root

OpenStudy (anonymous):

mathhelpplz your no help you have to solve this problem by calculus